Determine the equation of the line described. Put answer in the slope-intercept form, if possible.
Through (-2,-7) parallel to -6x - 5y = 57.
I tried to work through the problem and got the answer y = -6/5x - 23/5. Is that correct?
Thanks so much!!

Answer by vertciel(183)

Hi,
I recategorised your question to Linear Equations.
No, your answer is not correct. We can plug (-2,-7) in to check.
y = -6/5(-2) - 23/5
y = [-6)(-2) - 23]/5
= (12 - 23)/5
= -11/5
= -2.2 <-- Just a bit off -2.
Try again, and if you need more help, write back! Good luck!